# Dissipating quartets of excitations in a superconducting circuit

### v2

@@ I Introduction
Quantum Error Correction (QEC) relies on the measurement of error syndromes to detect spurious evolutions of a physical system embedding quantum information [1]. For stabilizer codes [2], the regular approach consists in measuring a set of stabilizers and decoding the measurement record to apply corrective feedback or feedforward controls. Alternatively, QEC can be made autonomous by engineering adequate dissipation channels [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. Autonomous QEC is only practical in the most simple decoding cases, in particular when each feedback control depends on the instantaneous value of a single stabilizer only. This is the case for most single-mode bosonic codes [22], in which a qubit encoded in an oscillator is stabilized by a few operators only. This approach was pioneered in the setting of two-legged cat qubits [10, 23, 24, 25, 26, 27, 28, 29]. The two-legged cat code manifold is spanned by two coherent states |±$|\pm\alpha\rangle$ of a target oscillator $a$. It is stabilized by a Lindblad dynamics entailed by a single dissipation channel $\dot{\bm{\rho}}_{a}=\gamma_{2}\mathcal{D}[{\bf a}^{2}-\alpha^{2}]\bm{\rho}_{a}$, where $\gamma_{2}$ is the dissipation rate and $\mathcal{D}[{\bf L}]\bm{\rho}={\bf L}\bm{\rho}{\bf L}^{\dagger}-\big{(}{\bf L}^{\dagger}{\bf L}\bm{\rho}+\bm{\rho}{\bf L}^{\dagger}{\bf L}\big{)}/2$. The Lindblad operator ${\bf a}^{2}-\alpha^{2}$ is of order two in the annihilation and creation operators $({\bf a},{\bf a^{\dagger}})$ and by extension, we characterize the stabilizing dissipation channel as second order. It may be engineered by activating a two-to-one photon exchange interaction between the target oscillator and an auxiliary mode $b$, combined with a resonant drive on $b$ to yield the Hamiltonian ${\bf H}_{2}=g_{2}({\bf a}^{2}-\alpha^{2}){\bf b^{\dagger}}+\mathrm{h.c.}$ If the auxiliary mode dissipates excitations into a cold environment at rate $\kappa_{b}$ much larger than the two-to-one photon exchange rate $g_{2}$, the multimode dynamics reduces to the desired single-mode Lindblad dynamics of the target mode with $\gamma_{2}=4g_{2}^{2}/\kappa_{b}$ [30, 31] [^1] This hierarchy is needed to adiabatically eliminate the mode $b$ from the dynamics, but does not need to be respected for the cat manifold to be stabilized [31]. Unless the two mode resonance frequencies are precisely targeted to reach $\omega_{b}=2\omega_{a}$ [28], the interaction needs to be parametrically modulated at $2\omega_{a}-\omega_{b}$ to become resonant. This can be realized with a variety of Josephson circuits such as the one pictured in Fig. 1a, all relying on the four-wave mixing capabilities of Josephson junctions. This strategy was successfully deployed to robustly suppress bit-flips of two-legged cat qubits. On the down side, this code does not offer any protection against phase-flips entailed by single-photon loss in the target mode, which is the dominant noise channel in most experimental platforms.
@@
Figure 1: Dissipation engineering in a Josephson circuit. a). A target oscillator (the memory, blue) is coupled to an auxiliary dissipative mode (the buffer, orange) by a SQUID participating in both modes. When the circuit loops are biased with magnetic fluxes $\Phi_{1}$ and $\Phi_{2}$ following Eq. (1) with $\epsilon(t)$ a pump oscillating at $4\omega_{a}-\omega_{b}$, the SQUID mediates a Kerr-free four-to-one photon exchange interaction, opening a four-photon dissipation channel acting on the memory. The mode impedances $Z_{a,b}$ set the strength of the interaction and are of the order of $R_{q}/10=650~{}\Omega$ in our experiment. b) Energy level-diagram featuring the four-to-one photon pump (purple wriggled arrow) and the buffer single-photon dissipation (orange wriggled arrow). In presence of strong non-linearity, energy levels are unevenly spaced and the pump drives a single transition only (dashed purple arrow represents an off-resonantly driven transition). c) Schematic representation of the circuit used in our experiment. A three-pad superconducting structure hosts the memory and buffer modes with respective anti-symmetric and symmetric voltages across the pads. A distributed $\lambda/2$ resonator capacitively coupled to the structure forms a third mode employed for readout. Foster synthesis of the linear circuit connected to the SQUID yields an effective circuit as in (a) but for the presence of a third mode and for series inductances in the arms of the SQUID, responsible for spurious non-linearities (see Fig. 2a and text). Magnetic biases are applied through two flux lines (top). A charge line (bottom) is capacitively connected to the circuit and enables linear drives and single-photon dissipation of the buffer. All lines are filtered by periodic loading to limit single-photon dissipation in the memory [40]. d) Optical microscope image of the sample, featuring all the structures pictured in (c) with matching false colors, surrounded by a ground plane (brown).
@@
where $\varphi_{0}=\hbar/(2e)$ is the reduced flux quantum and the AC signal $\epsilon$ is referred to as the pump. The SQUID couples to the sum of the modes reduced phase operators $\bm{\varphi}=\varphi^{ZPF}_{a}({\bf a}+{\bf a}^{\dagger})+\varphi^{ZPF}_{b}({\bf b}+{\bf b}^{\dagger})$ via
${\bf H}_{J}(t)$ $=-E_{J}\cos(\bm{\varphi}-\frac{\Phi_{1}}{\varphi_{0}})-E_{J}\cos(\bm{\varphi}+\frac{\Phi_{2}}{\varphi_{0}})$ (2)
$=2E_{J}~{}\sin\big{(}\epsilon(t)\big{)}\sin(\bm{\varphi})$
@@ II.1 Design
To model and quantize our system, we perform Foster synthesis of the linear circuit connected to the SQUID [44, 41, 43], yielding the three system modes (labeled $a$, $b$, $r$) represented by a chain of parallel $LC$ oscillators with the hierarchy $Z_{r}\ll Z_{a}\sim Z_{b}\sim R_{q}/10$ (see Fig. 2a). The ATS shunt inductor is accounted for in this synthesis, but not the SQUID itself, whose arms each feature a junction and a small series inductance. The SQUID is subsequently connected in parallel with the chain to obtain the effective circuit pictured in Fig. 2a. It is analogous to the ideal circuit pictured in Fig. 1a but for the presence of a third mode and for the series inductances in the arms of the SQUID. The fluxes (1) biasing the loops of the ATS are delivered through two dedicated feedlines, denoted as flux lines 222 The fluxes $\Phi_{1}$ and $\Phi_{2}$ that thread the physical circuit loops can be formally allocated to each arm of the SQUID, such that they are the same fluxes threading the effective circuit loops.. A third feedline, denoted as the charge line, is capacitively coupled to the three-pad structure and to the readout resonator. It allows direct linear drive of the modes and sets the single-photon decay rate of the buffer and readout modes (see Table 1). Note that in principle, one could leverage the distinct symmetries of the memory and buffer modes to selectively couple a feedline to the buffer, thus preventing radiative decay of memory photons. This was not attempted in our experiment, in which the coupling of the feedline is of the same order for all modes. Instead, we filter all feedlines by sinusoidally modulating their geometry, thereby suppressing radiative decay at the memory frequency [40].
@@ II.2 Modes dispersion against bias fluxes
In order to characterize the system, we record the modes resonance frequencies against DC magnetic fluxes biasing the ATS loops in absence of pump ($\epsilon(t)=0$). We first probe the readout mode in single-tone spectroscopy through the charge line, yielding the frequency map represented in Fig. 2b. We identify the ATS working point corresponding to the DC biases introduced in Eq. (1) as a saddle point of the frequency map, along with analog points located at half-integer numbers of flux quanta away. This allows us to relate the direct currents applied through the flux lines to the actual fluxes threading the ATS (Fig. 2b uses calibrated flux coordinates  [^3] The fluxes $\Phi_{1}$ and $\Phi_{2}$ are known up to an integer number of flux quanta and depend on applied currents through an affine transformation. The map is centered at non-zero applied currents, indicating significant flux offsets., and two non-equivalent saddle points are identified by crossing arrows). Next, we record the memory and buffer resonance frequencies—respectively in two-tone spectroscopy leveraging the memory-readout dispersive coupling and in single-tone spectroscopy—around the saddle points. Fig. 2c features each mode resonance frequency (empty circles) when sweeping applied fluxes through the saddle points of the frequency map, along directions of maximum and minimum curvatures (indicated by double-headed arrows with matching colors in Fig. 2b). Strikingly, the modes have near identical resonance frequencies at both saddle points (markers in bright colors corresponding to the saddle point $\mathrm{SP}_{1}$ are aligned with markers in dim colors corresponding to the saddle point $\mathrm{SP}_{2}$), indicating that the Josephson junctions forming the SQUID have a relative energy difference smaller than $0.01~{}\%$ [40]. We attribute this exceptionally low circuit asymmetry to chance in fabrication and junction aging behavior. Indeed, the asymmetry was found to be two orders of magnitude higher in similar circuits that we have characterized, and an order of magnitude higher during a previous cooldown of the same circuit.
@@ II.3 Stray inductances and memory non-linearity
Unexpectedly, when driving the memory on resonance and subsequently probing the readout resonator, we detect a reflected signal that oscillates with the duration and amplitude of the memory drive pulse [40]. Measuring such Rabi oscillations unambiguously reveals a significant non-linearity of the memory, whose transitions can be addressed individually. Driving the memory with sequences of charge drives (to excite single-photon transitions $|n\rangle_{a}\leftrightarrow|n+1\rangle_{a}$) and/or flux drives (to excite three-photon transitions $|n\rangle_{a}\leftrightarrow|n+3\rangle_{a}$ leveraging the wave-mixing properties of the ATS), we reconstruct the memory mode energy level structure at the saddle point (see Fig. 2d). The anharmonicity $\omega_{a}^{n+1,n+2}-\omega_{a}^{n,n+1}$ is positive for all $n$, lies in the 1—10 MHz range and decreases sharply with $n$. As discussed in Sec. IV, we unambiguously relate this exotic pattern to a Cooper-pair pairing mechanism (renormalization of the Josephson potential with an added $\cos(2\varphi)$ contribution) owing to stray inductance in the tracks of the SQUID [47, 40] (inductive energy $\epsilon_{L}\gg E_{J}$, see Fig. 2a). Including this mechanism in our model, we quantitatively reproduce all data presented in Fig. 2. The fitted parameters, presented in Table 1, are reasonably close to design values.
@@
Table 1: Fitted circuit parameters. $\delta E_{J}$ is the difference in Josephson energy between the two junctions of the SQUID, while $E_{J}$ labels the mean energy. Details of the fit procedure are presented as Supplemental materials.
@@ III.1 Preparing a Fock state and measuring its occupation
In order to characterize the engineered dissipation channel, we need to prepare the memory in each Fock state $|n\rangle_{a}$  [^4] The spurious non-linear term in the circuit Hamiltonian is a small perturbation so that eigenstates have an overlap close to 1 with unperturbed Fock states. and to monitor the time decay of its population. Both the preparation and measurement are enabled by the same non-linearity preventing us from addressing more than a single four-to-one photon transition. In principle, the former only requires to apply sequences of $\pi$-pulses on the $|k\rangle_{a}\leftrightarrow|k+1\rangle_{a}$ transitions for $k<n$, while the latter is enabled by the dispersive shift of the readout mode resonance frequency conditioned on $n$. However, preparation and measurement of highly excited states are challenging due to the memory anharmonicity and the readout mode dispersive shift both tailing off with the level index $n$. We therefore limit ourselves to preparation and measurement of Fock states up to $n=4$.
@@
In detail, the states are prepared with pulse sequences combining $\pi$-pulses applied on the single-photon transitions $|0\rangle_{a}\leftrightarrow|1\rangle_{a}$ and $|1\rangle_{a}\leftrightarrow|2\rangle_{a}$ (direct charge drive) and/or on the three-photon transitions $|0\rangle_{a}\leftrightarrow|3\rangle_{a}$ and $|1\rangle_{a}\leftrightarrow|4\rangle_{a}$ (flux drive of the ATS). The duration of these pulses being on par with the coherence time of the transitions, the preparation is imperfect. It is nonetheless sufficient to estimate the lifetime of each state, as detailed in the next section. As for the readout, the ground state can be directly distinguished from excited states through the pull exerted by excited states on the readout mode resonance frequency. However, their respective pulls saturate rapidly with increasing state index, making them hard to distinguish from one another. We circumvent this shortcoming with the scheme depicted in Fig. 3a. To estimate the occupation of state $|n\rangle_{a}$ ($n\geq 1$), we apply or not a $\pi$-pulse on the $|n\rangle_{a}\leftrightarrow|n+3\rangle_{a}$ transition before probing the readout mode on resonance with a long pulse. Since the dynamics of the decay from $|n+3\rangle_{a}$ to the ground state is delayed compared to the decay from $|n\rangle_{a}$, we detect a time-shifted signal in the reflected field $s_{n}^{ON}(t)$ detected after the application of the $\pi$-pulse with respect to the field $s_{n}^{OFF}(t)$ detected in absence of the pulse [40]. We integrate the detected fields with a time-envelope $\mathcal{E}(t)$ chosen to maximize the signal-to-noise ratio of the differential signal $S_{n}=\int_{t}\Re\Big{(}\big{(}s_{n}^{ON}(t)-s_{n}^{OFF}(t)\big{)}\mathcal{E}(t)\Big{)}\mathrm{d}t$ between these two situations [40]. Since the $\pi$-pulse only modifies the memory state when $|n\rangle_{a}$ is occupied, the integrated signal $S_{n}$ is proportional to the occupation of state $|n\rangle_{a}$ at the instant when the pulse is delivered, even if population transfer is imperfect. Note that both for state preparation and detection, we favor pulses applied to three-photon transitions to ones applied to single-photon transitions as the detuning between neighboring three-photon transitions is larger, allowing for stronger drives.
@@
Before applying these preparation and measurement methods to characterize the memory population dynamics in presence of a pump, we report that some of the memory transitions—driven either during preparation or measurement, involving a single photon or three—were frequently and transiently observed to shift or split [40]. We attribute this effect to hybridization of the memory to a bath of Two-Level-Systems (TLS). Such TLS baths, whose frequency landscape reconfigure on timescales of the order of a second, are frequently observed in superconducting circuits [49, 50]. While the exact coupling mechanism and density of TLS’s was not characterized in our experiment, at least one of the memory transitions was typically affected by a TLS at any given time so that waiting for the bath to reconfigure in a more favorable configuration (either passively or cycling the system to room temperature) was not a viable option. Instead, we use an off-resonant flux drive of large amplitude to shift the frequency of the memory transitions away from any resonant coupling with TLS’s [40]. The frequency of this drive was chosen to minimize decoherence induced on the memory, but still leads to degradation of its energy relaxation and dephasing times 555 Note that in principle, the memory coherence should not be affected by this drive as the ATS remains at a dynamical sweet-spot, in the sense that its susceptibility to magnetic flux drifts averages out over a period of the drive. The datasets presented in the following section were acquired in presence of such a drive.

### v3

@@ I Introduction
Quantum Error Correction (QEC) relies on the measurement of error syndromes to detect spurious evolutions of a physical system embedding quantum information [42]. For stabilizer codes [17], the regular approach consists in measuring a set of stabilizers and decoding the measurement record to apply corrective feedback or feedforward controls. Alternatively, QEC can be made autonomous by engineering adequate dissipation channels [84, 87, 30, 3, 62, 76, 47, 36, 8, 1, 22, 23, 57, 31, 61, 27, 10, 85, 59]. Autonomous QEC is only practical in the most simple decoding cases, in particular when each feedback control depends on the instantaneous value of a single stabilizer only. This is the case for most single-mode bosonic codes [2], in which a qubit encoded in an oscillator is stabilized by a few operators only. This approach was pioneered in the setting of two-legged cat qubits [36, 28, 78, 29, 4, 55, 34, 53]. The two-legged cat code manifold is spanned by two coherent states $|\pm\alpha\rangle$ of a target oscillator $a$. It is stabilized by a Lindblad dynamics entailed by a single dissipation channel $\dot{\boldsymbol{\rho}}_{a}=\gamma_{2}\mathcal{D}[{\bf a}^{2}-\alpha^{2}]\boldsymbol{\rho}_{a}$, where $\gamma_{2}$ is the dissipation rate and $\mathcal{D}[{\bf L}]\boldsymbol{\rho}={\bf L}\boldsymbol{\rho}{\bf L}^{\dagger}-\big({\bf L}^{\dagger}{\bf L}\boldsymbol{\rho}+\boldsymbol{\rho}{\bf L}^{\dagger}{\bf L}\big)/2$. The Lindblad operator ${\bf a}^{2}-\alpha^{2}$ is of order two in the annihilation and creation operators $({\bf a},{\bf a^{\dagger}})$ and by extension, we characterize the stabilizing dissipation channel as second order. It may be engineered by activating a two-to-one photon exchange interaction between the target oscillator and an auxiliary mode $b$, combined with a resonant drive on $b$ to yield the Hamiltonian ${\bf H}_{2}=g_{2}({\bf a}^{2}-\alpha^{2}){\bf b^{\dagger}}+\mathrm{h.c.}$ If the auxiliary mode dissipates excitations into a cold environment at rate $\kappa_{b}$ much larger than the two-to-one photon exchange rate $g_{2}$, the multimode dynamics reduces to the desired single-mode Lindblad dynamics of the target mode with $\gamma_{2}=4g_{2}^{2}/\kappa_{b}$ [37, 58, 74]. Unless the two mode resonance frequencies are precisely targeted to reach $\omega_{b}=2\omega_{a}$ [34], the interaction needs to be parametrically modulated at $2\omega_{a}-\omega_{b}$ to become resonant. This can be realized with a variety of Josephson circuits such as the one pictured in Fig. 1a, all relying on the four-wave mixing capabilities of Josephson junctions. This strategy was successfully deployed to robustly suppress bit-flips of two-legged cat qubits. On the down side, this code does not offer any protection against phase-flips entailed by single-photon loss in the target mode, which is the dominant noise channel in most experimental platforms.
@@
Figure 1: Dissipation engineering in a Josephson circuit. a). A target oscillator (the memory, blue) is coupled to an auxiliary dissipative mode (the buffer, orange) by a SQUID participating in both modes. When the circuit loops are biased with magnetic fluxes $\Phi_{1}$ and $\Phi_{2}$ following Eq. (1) with $\epsilon(t)$ a pump oscillating at $4\omega_{a}-\omega_{b}$, the SQUID mediates a Kerr-free four-to-one photon exchange interaction, opening a four-photon dissipation channel acting on the memory. The mode impedances $Z_{a,b}$ set the strength of the interaction and are of the order of $R_{q}/10=650~\Omega$ in our experiment. b) Energy level-diagram featuring the four-to-one photon pump (purple wriggled arrow) and the buffer single-photon dissipation (orange wriggled arrow). In presence of strong non-linearity, energy levels are unevenly spaced and the pump drives a single transition only (dashed purple arrow represents an off-resonantly driven transition). c) Schematic representation of the circuit used in our experiment. A three-pad superconducting structure hosts the memory and buffer modes with respective anti-symmetric and symmetric voltages across the pads. A distributed $\lambda/2$ resonator capacitively coupled to the structure forms a third mode employed for readout. Foster synthesis of the linear circuit connected to the SQUID yields an effective circuit as in (a) but for the presence of a third mode and for series inductances in the arms of the SQUID, responsible for spurious non-linearities (see Fig. 2a and text). Magnetic biases are applied through two flux lines (top). A charge line (bottom) is capacitively connected to the circuit and enables linear drives and single-photon dissipation of the buffer. All lines are filtered by periodic loading to limit single-photon dissipation in the memory (see Sec. A.3). d) Optical microscope image of the sample, featuring all the structures pictured in (c) with matching false colors, surrounded by a ground plane (brown).
@@
where $\varphi_{0}=\hbar/(2e)$ is the reduced flux quantum and the AC signal $\epsilon$ is referred to as the pump. The SQUID couples to the sum of the modes reduced phase operators $\boldsymbol{\varphi}=\varphi^{ZPF}_{a}({\bf a}+{\bf a}^{\dagger})+\varphi^{ZPF}_{b}({\bf b}+{\bf b}^{\dagger})$ via
${\bf H}_{J}(t)$ $=-E_{J}\cos(\boldsymbol{\varphi}-\frac{\Phi_{1}}{\varphi_{0}})-E_{J}\cos(\boldsymbol{\varphi}+\frac{\Phi_{2}}{\varphi_{0}})$ (2)
$=2E_{J}~\sin\big(\epsilon(t)\big)\sin(\boldsymbol{\varphi})$
@@ II.1 Design
To model and quantize our system, we perform Foster synthesis of the linear circuit connected to the SQUID [12, 43, 35], yielding the three system modes (labeled $a$, $b$, $r$) represented by a chain of parallel $LC$ oscillators with the hierarchy $Z_{r}\ll Z_{a}\sim Z_{b}\sim R_{q}/10$ (see Fig. 2a). The ATS shunt inductor is accounted for in this synthesis, but not the SQUID itself, whose arms each feature a junction and a small series inductance. The SQUID is subsequently connected in parallel with the chain to obtain the effective circuit pictured in Fig. 2a. It is analogous to the ideal circuit pictured in Fig. 1a but for the presence of a third mode and for the series inductances in the arms of the SQUID. The fluxes (1) biasing the loops of the ATS are delivered through two dedicated feedlines, denoted as flux lines [70]. A third feedline, denoted as the charge line, is capacitively coupled to the three-pad structure and to the readout resonator. It allows direct linear drive of the modes and sets the single-photon decay rate of the buffer and readout modes (see Table 1). Note that in principle, one could leverage the distinct symmetries of the memory and buffer modes to selectively couple a feedline to the buffer, thus preventing radiative decay of memory photons. This was not attempted in our experiment, in which the coupling of the feedline is of the same order for all modes. Instead, we filter all feedlines by sinusoidally modulating their geometry as detailed in Sec. A.3, thereby suppressing radiative decay at the memory frequency.
@@ II.2 Modes dispersion against bias fluxes
In order to characterize the system, we record the modes resonance frequencies against DC magnetic fluxes biasing the ATS loops in absence of pump ($\epsilon(t)=0$). We first probe the readout mode in single-tone spectroscopy through the charge line, yielding the frequency map represented in Fig. 2b. We identify the ATS working point corresponding to the DC biases introduced in Eq. (1) as a saddle point of the frequency map, along with analog points located at half-integer numbers of flux quanta away. This allows us to relate the direct currents applied through the flux lines to the actual fluxes threading the ATS (Fig. 2b uses calibrated flux coordinates [69], and two non-equivalent saddle points are identified by crossing arrows). Next, we record the memory and buffer resonance frequencies—respectively in two-tone spectroscopy leveraging the memory-readout dispersive coupling and in single-tone spectroscopy—around the saddle points. Fig. 2c features each mode resonance frequency (empty circles) when sweeping applied fluxes through the saddle points of the frequency map, along directions of maximum and minimum curvatures (indicated by double-headed arrows with matching colors in Fig. 2b). Strikingly, the modes have near identical resonance frequencies at both saddle points (markers in bright colors corresponding to the saddle point $\mathrm{SP}_{1}$ are aligned with markers in dim colors corresponding to the saddle point $\mathrm{SP}_{2}$), indicating that the Josephson junctions forming the SQUID have a relative energy difference smaller than $0.01~\%$ (see Sec. D.1). We attribute this exceptionally low circuit asymmetry to chance in fabrication and junction aging behavior. Indeed, the asymmetry was found to be two orders of magnitude higher in similar circuits that we have characterized, and an order of magnitude higher during a previous cooldown of the same circuit.
@@ II.3 Stray inductances and memory non-linearity
Unexpectedly, when driving the memory on resonance and subsequently probing the readout resonator, we detect a reflected signal that oscillates with the duration and amplitude of the memory drive pulse. Measuring such Rabi oscillations unambiguously reveals a significant non-linearity of the memory, whose transitions can be addressed individually. Driving the memory with sequences of charge drives (to excite single-photon transitions $|n\rangle_{a}\leftrightarrow|n+1\rangle_{a}$) and/or flux drives (to excite three-photon transitions $|n\rangle_{a}\leftrightarrow|n+3\rangle_{a}$ leveraging the wave-mixing properties of the ATS), we reconstruct the memory mode energy level structure at the saddle point (see Fig. 2d). The anharmonicity $\omega_{a}^{n+1,n+2}-\omega_{a}^{n,n+1}$ is positive for all $n$, lies in the 1—10 MHz range and decreases sharply with $n$. As discussed in Sec. IV, we unambiguously relate this exotic pattern to a Cooper-pair pairing mechanism (renormalization of the Josephson potential with an added $\cos(2\varphi)$ contribution) owing to stray inductance in the tracks of the SQUID [65] (inductive energy $\epsilon_{L}\gg E_{J}$, see Fig. 2a). Including this mechanism in our model, we quantitatively reproduce all data presented in Fig. 2. The fitted parameters, presented in Table 1, are in good agreement with design values (see Sec. B for details).
@@
Table 1: Fitted circuit parameters. $\delta E_{J}$ is the difference in Josephson energy between the two junctions of the SQUID, while $E_{J}$ labels the mean energy. Details of the fit procedure are presented in Sec. D.
@@ III.1 Preparing a Fock state and measuring its occupation
In order to characterize the engineered dissipation channel, we need to prepare the memory in each Fock state $|n\rangle_{a}$ [72] and to monitor the time decay of its population. Both the preparation and measurement are enabled by the same non-linearity preventing us from addressing more than a single four-to-one photon transition. In principle, the former only requires to apply sequences of $\pi$-pulses on the $|k\rangle_{a}\leftrightarrow|k+1\rangle_{a}$ transitions for $k<n$, while the latter is enabled by the dispersive shift of the readout mode resonance frequency conditioned on $n$. However, preparation and measurement of highly excited states are challenging due to the memory anharmonicity and the readout mode dispersive shift both tailing off with the level index $n$. We therefore limit ourselves to preparation and measurement of Fock states up to $n=4$.
@@
In detail, the states are prepared with pulse sequences combining $\pi$-pulses applied on the single-photon transitions $|0\rangle_{a}\leftrightarrow|1\rangle_{a}$ and $|1\rangle_{a}\leftrightarrow|2\rangle_{a}$ (direct charge drive) and/or on the three-photon transitions $|0\rangle_{a}\leftrightarrow|3\rangle_{a}$ and $|1\rangle_{a}\leftrightarrow|4\rangle_{a}$ (flux drive of the ATS). The duration of these pulses being on par with the coherence time of the transitions, the preparation is imperfect. It is nonetheless sufficient to estimate the lifetime of each state, as detailed in the next section. As for the readout, the ground state can be directly distinguished from excited states through the pull exerted by excited states on the readout mode resonance frequency. However, their respective pulls saturate rapidly with increasing state index, making them hard to distinguish from one another. We circumvent this shortcoming with the scheme depicted in Fig. 3a. To estimate the occupation of state $|n\rangle_{a}$ ($n\geq 1$), we apply or not a $\pi$-pulse on the $|n\rangle_{a}\leftrightarrow|n+3\rangle_{a}$ transition before probing the readout mode on resonance with a long pulse. Since the dynamics of the decay from $|n+3\rangle_{a}$ to the ground state is delayed compared to the decay from $|n\rangle_{a}$, we detect a time-shifted signal in the reflected field $s_{n}^{ON}(t)$ detected after the application of the $\pi$-pulse with respect to the field $s_{n}^{OFF}(t)$ detected in absence of the pulse (see Sec. C.1). We integrate the detected fields with a time-envelope $\mathcal{E}(t)$ chosen to maximize the signal-to-noise ratio of the differential signal $S_{n}=\int_{t}\Re\Big(\big(s_{n}^{ON}(t)-s_{n}^{OFF}(t)\big)\mathcal{E}(t)\Big)\mathrm{d}t$ between these two situations. Since the $\pi$-pulse only modifies the memory state when $|n\rangle_{a}$ is occupied, the integrated signal $S_{n}$ is proportional to the occupation of state $|n\rangle_{a}$ at the instant when the pulse is delivered, even if population transfer is imperfect. Note that both for state preparation and detection, we favor pulses applied to three-photon transitions to ones applied to single-photon transitions as the detuning between neighboring three-photon transitions is larger, allowing for stronger drives.
@@
Before applying these preparation and measurement methods to characterize the memory population dynamics in presence of a pump, we report that some of the memory transitions—driven either during preparation or measurement, involving a single photon or three—were frequently and transiently observed to shift or split (see Sec. E). We attribute this effect to hybridization of the memory to a bath of Two-Level-Systems (TLS). Such TLS baths, whose frequency landscape reconfigure on timescales of the order of a second, are frequently observed in superconducting circuits [46, 38]. While the exact coupling mechanism and density of TLS’s was not characterized in our experiment, at least one of the memory transitions was typically affected by a TLS at any given time so that waiting for the bath to reconfigure in a more favorable configuration (either passively or cycling the system to room temperature) was not a viable option. Instead, we use an off-resonant flux drive of large amplitude to shift the frequency of the memory transitions away from any resonant coupling with TLS’s. The frequency of this drive was chosen to minimize decoherence induced on the memory, but still leads to degradation of its energy relaxation and dephasing times [44]. The datasets presented in the following section were acquired in presence of such a drive.